# Optimal Stopping with a Predicted Prior

### v1

@@ A.2 Asymptotic Characterization of $\theta_{n}^{*}$
Recall Equation 5:
$\int_{s}^{1}\frac{\big(\frac{1-t}{\theta^{*}_{n}(s)}+t\big)^{n-1}-1}{1-t}\,\mathrm{d}t=1~.$
Fix any $s$, substitute $x_{n}=\frac{1}{\theta^{*}_{n}(s)}-1$, the left-hand side becomes
$\int_{s}^{1}\frac{\left(1+x_{n}(1-t)\right)^{n-1}-1}{1-t}\,\mathrm{d}t~.$
Expanding the left-hand side with the binomial theorem, we have
$(n-1)(1-s)x_{n}+\binom{n-1}{2}\frac{(1-s)^{2}}{2}x_{n}^{2}+\cdots+\binom{n-1}{k}\frac{(1-s)^{k}}{k!}x_{n}^{k}+\cdots+\frac{(1-s)^{n-1}}{(n-1)!}x_{n}^{n-1}=1~.$
Intuitively, as $n\to\infty$, the coefficients $\frac{(n-2)(n-3)\cdots(n-k)}{(n-1)^{k-1}}\to 1$ for any fixed $k$. Thus, we expect that $c$ satisfies
$\sum_{k\geq 1}\frac{c^{k}}{k!k}=1~,$
which defines the constant $\gamma\approx 0.80435$ in the lemma.
Consider the difference:
It follows that the difference is bounded by
Finally, we conclude that $(n-1)(1-s)x_{n}\to\gamma$ as $n\to\infty$, which implies that $x_{n}=\frac{\gamma}{(n-1)(1-s)}+o(\tfrac{1}{n})$. This completes the proof by $\theta^{*}_{n}(s)=\frac{1}{1+x_{n}}=\frac{1}{1+\frac{\gamma}{(n-1)(1-s)}}+o(\tfrac{1}{n})$. ∎
@@ Appendix B Numerical Solution to the Differential Equation in Theorem 1
We assume throughout that $\theta$ is left‑continuous, non‑increasing, and $\theta(s)\equiv 0$ for $s\in[\lambda_{2},1]$.
@@
Here, the last equality is obtained by exchanging the order of integration and using $\theta\equiv 0$ on $\lambda_{2},1]$.
@@
where the left-hand side decreases from $\infty$ to 0 as $\theta_{m}$ increases from 0 to $\infty$. Thus, there exists a $\theta_{m}^{*}>0$ such that the equality holds. [^9] During experiment, we use binary search to locate $\theta_{m}^{*}$ with $100$ iterations so that the error can be ignored.
@@
By unifying intervals that $\theta_{i}$ exceeds $1$, as Lemma 22, we only need to check that Equation 21 holds for $\min\{\theta_{i},1\}$, where $i$ is the last $\theta_{i}$ exceeding $1$. This instantly stands by the non-increasing of the left-hand side. ∎
@@
If an $\alpha$ such that the corresponding recursively constructed sequence $(\theta_{i}^{*})_{i=1}^{m}$ satisfies $\theta_{1}^{*}\geq 1$, then the step function $\theta^{*}$ with step values $(\min\{\theta_{i}^{*},1\})_{i=1}^{m}$ is a left-continuous non-increasing function satisfying Equation 17 and Equation 18.
@@
Intuitively, as $m\to\infty$, the step function approximation tightens. In experiments, we set $m=300$. For details, see the implementation of the backward recursion as well as the solved $\theta_{i}^{*}$’s in our code repository.
Figure 3 illustrates the case when $\beta=0.01$ and $m=300$ yields $\alpha\approx 0.6908$ with $\theta^{*}_{1}\approx 1.0165$. This implies that $\mathcal{A}(F,(\theta^{*})^{\nicefrac{{1}}{{n}}})$ with $\theta^{*}$ being the step function with step values $(\min\{\theta_{i}^{*},1\})_{i=1}^{300}$ is an $\alpha\approx 0.6908$-consistent and $\beta=0.01$-robust algorithm.
@@
Appendix C Missing Proofs for MaxProb Hardness
C.1 Minor-Oblivious Algorithms: Proof of Proposition 17
The key observation is that, when the current value is best-so-far and equal across sequences, the conditional probability of stopping at time $t$ is invariant across all truncation levels $k\geq\ell$ for which the conditioning event has positive probability. This is formalized in the following lemma.
@@ C.1 Minor-Oblivious Algorithms: Proof of Proposition 17
That is, the probability of accepting a best-so-far value $x_{t}=\ell$ equals the probability that $\mathcal{A}$ accepts at time $t$, conditioned on 1) $x_{t}=\ell$; 2) $x_{t}$ is best-so-far; and 3) $\mathcal{A}$ rejects the first t1t-1 values. By Lemma 26, we may define this conditional probability with respect to prior $F_{K}$ without loss of generality.
@@
C.2 Relating Rejection and Acceptance Probabilities: Proof of Lemma 18

### v2

@@ A.1 MaxExp: Optimal Threshold
With $g(u,t)$ and $h(t,x)$ defined above, Lemma 22 implies that for each fixed $t\in[0,1)$ the function $h(t,\cdot)$ is strictly increasing on $0,\infty)$ and has a unique zero at $x=x^{*}(t)$.
Fix $x\geq 0$ and let $u=F(x)$. For $u\in[0,1)$ and $t\in[0,1)$,
$\partial_{t}g(u,t)=-\frac{(n-1)u^{n-1}}{(1-t+tu)^{n}}\leq 0,$
so $g(F(x),t)$ is non-increasing in $t$, $1/g(F(x),t)$ is non-decreasing, and hence
$\partial_{t}h(t,x)=x\,\partial_{t}\left[\frac{1}{g(F(x),t)}\right]\geq 0.$
Thus, for any $0\leq t_{1}<t_{2}\leq 1$ and any $x\geq 0$,
$h(t_{1},x)\leq h(t_{2},x).$
Fix $0\leq t_{1}<t_{2}\leq 1$. Since $x^{*}(t_{2})$ is the unique zero of $h(t_{2},\cdot)$, we have $h(t_{2},x^{*}(t_{2}))=0$, and the inequality above gives $h(t_{1},x^{*}(t_{2}))\leq 0$. Because $h(t_{1},\cdot)$ is strictly increasing on $0,\infty)$ and has a unique zero at $x^{*}(t_{1})$, this implies $x^{*}(t_{1})\geq x^{*}(t_{2})$.
Hence $x^{*}(t)$ is non-increasing in $t$ on $[0,1]$. ∎
A.2 MaxProb: Optimal Threshold
Lemma 24 (Optimal threshold function for MaxProb).
For any $n\in\mathbb{N}$ and any prior $F$, the optimal threshold function $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}$ that maximizes the consistency of $\mathcal{A}(F,\theta_{n}^{\scriptscriptstyle\mathrm{prob}})$ under MaxProb is uniquely determined at each time $t$ by the unique solution $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t)\in[0,1]$ of
$\psi(t,q)~=~0,$ (19)
where
$\psi(t,q)~\coloneqq~\int_{t}^{1}\frac{(1-s+sq)^{n-1}-q^{n-1}}{1-s}\,\,\mathrm{d}s-q^{n-1}.$
Moreover, $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t)$ is non-increasing in $t$ on $[0,1]$.
Note that the form of Equation (19) coincides with the equation defining the threshold function used by Example 2.
Proof.
Fix $t\in[0,1]$ and suppose the thresholds $\theta(s)$ have already been fixed for all $s\in(t,1]$. Let $x$ denote the value that arrives at time $t$, and assume $x$ is a record (otherwise the algorithm cannot accept at $t$). By definition, the algorithm $\mathcal{A}(F,\theta)$ accepts this value $x$ at time $t$ if and only if $F(x)>\theta(t)$. By optimality, the threshold chosen at time $t$ must make the success probabilities from accepting and rejecting this value $x$ equal.
Accepting. If the algorithm accepts a record $x$ at time $t$, the success event is that $x$ is the largest among all $n$ values. Conditional on $x$ being a record at time $t$, this occurs with probability
[…]
Rejecting. If instead the algorithm rejects such an $x$, then by the non-increasing threshold assumption it will accept the next value larger than $x$, which is necessarily the next record that exceeds the threshold. Given that there are in total $k$ values larger than $x$, the probability that the algorithm eventually picks the largest among all values is $1/k$, since the arrival order is uniform at random and the algorithm accepts the first exceedance over $x$. The probability that there are exactly $k$ values larger than $x$ and arriving after time $t$, conditional on $x$ being a record at time $t$, is
$\frac{\binom{n-1}{k}\big((1-t)(1-F(x))\big)^{k}F(x)^{n-1-k}}{(1-t+tF(x))^{n-1}}.$
Equating the success probability from accepting and rejecting $x$ yields the backward-induction equation
$\sum_{k=1}^{n-1}\binom{n-1}{k}\frac{1}{k}\big((1-t)(1-F(x))\big)^{k}F(x)^{n-1-k}~=~F(x)^{n-1}.$ (20)
For the quantile threshold $q=F(x)$, this becomes
$\sum_{k=1}^{n-1}\binom{n-1}{k}\frac{1}{k}\big((1-t)(1-q)\big)^{k}q^{n-1-k}~=~q^{n-1}.$
Using the identity $\frac{1}{k}=\int_{0}^{1}y^{k-1}\,\,\mathrm{d}y$, the left-hand side can be written as
$\sum_{k=1}^{n-1}\binom{n-1}{k}\frac{1}{k}\big((1-t)(1-q)\big)^{k}q^{n-1-k}$ $=\sum_{k=1}^{n-1}\binom{n-1}{k}q^{n-1-k}\int_{0}^{(1-t)(1-q)}y^{k-1}\,\,\mathrm{d}y$
$=\int_{0}^{(1-t)(1-q)}\frac{1}{y}\sum_{k=1}^{n-1}\binom{n-1}{k}q^{n-1-k}y^{k}\,\,\mathrm{d}y$
$=\int_{0}^{(1-t)(1-q)}\frac{(y+q)^{n-1}-q^{n-1}}{y}\,\,\mathrm{d}y$
$=\int_{t}^{1}\frac{(1-s+sq)^{n-1}-q^{n-1}}{1-s}\,\,\mathrm{d}s,$
where the last equality follows from the change of variables y=(1s)(1q)y=(1-s)(1-q). Under this substitution, $y=0$ corresponds to $s=1$ and y=(1t)(1q)y=(1-t)(1-q) corresponds to $s=t$, which reverses the limits. Thus, Equation 20 is equivalent to
$\int_{t}^{1}\frac{(1-s+sq)^{n-1}-q^{n-1}}{1-s}\,\,\mathrm{d}s~=~q^{n-1},$
which is precisely the equation $\psi(t,q)=0$ in Equation 19.
We now prove existence and uniqueness of the solution of $\psi(t,q)=0$ for each fixed $t$, and then establish the monotonicity of $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t)$.
For each fixed $t\in[0,1)$, the function $\psi(t,q)$ is continuous on $q\in[0,1]$. At $q=0$, the integrand equals $(1-s)^{n-2}$, hence
$\psi(t,0)=\int_{t}^{1}(1-s)^{n-2}\,\,\mathrm{d}s>0.$
At $q=1$, the integrand is identically zero, so
$\psi(t,1)=0-1^{n-1}=-1<0.$
Thus, for each $t\in[0,1)$ the equation $\psi(t,q)=0$ has at least one root $q\in(0,1)$.
For uniqueness, observe first that when $t=1$, the integrand in $\psi(1,q)$ is zero for all $q\in[0,1]$, so $\psi(1,q)=-q^{n-1}$ and $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(1)=0$ is the unique root. For $t<1$, define
$\tilde{\psi}(t,q)\coloneqq q^{-(n-1)}\psi(t,q)=\int_{t}^{1}\frac{\left(\frac{1-s+sq}{q}\right)^{n-1}-1}{1-s}\,\,\mathrm{d}s-1,\qquad q\in(0,1).$
Differentiating with respect to $q$ gives
$\partial_{q}\tilde{\psi}(t,q)=-\frac{n-1}{q^{n}}\int_{t}^{1}(1-s+sq)^{n-2}\,\,\mathrm{d}s<0,$
so for each fixed $t<1$ the function $\tilde{\psi}(t,q)$ is strictly decreasing in $q$ on $(0,1)$. Thus, for each $t\in[0,1)$ the equation $\tilde{\psi}(t,q)=0$ has a unique solution $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t)\in(0,1)$. Because $\tilde{\psi}(t,q)=0$ is equivalent to $\psi(t,q)=0$ for $q>0$, the solution of $\psi(t,q)=0$ is also unique.
Finally, it remains to show that $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t)$ is non-increasing in $t$. Fix $0\leq t_{1}<t_{2}<1$ and suppose, for contradiction, that $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t_{2})>\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t_{1})$. Since $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t_{i})$ is the unique zero of $\tilde{\psi}(t_{i},q)$, we have
$\tilde{\psi}(t_{i},\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t_{i}))=0,\qquad i=1,2.$
For fixed $t_{1}$, the function $q\mapsto\tilde{\psi}(t_{1},q)$ is strictly decreasing on $(0,1)$ and has a unique zero at $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t_{1})$. The inequality $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t_{2})>\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t_{1})$ therefore implies
$\tilde{\psi}(t_{1},\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t_{2}))<\tilde{\psi}(t_{1},\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t_{1}))=0.$ (21)
On the other hand, for fixed $q=\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t_{2})$ the function $\tilde{\psi}(t,q)$ is strictly decreasing on $[0,1)$. Since $t_{1}<t_{2}$, this yields
$\tilde{\psi}(t_{1},\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t_{2}))>\tilde{\psi}(t_{2},\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t_{2}))=0.$ (22)
The inequalities (21) and (22) contradict each other. Hence the assumption $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t_{2})>\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t_{1})$ is impossible, and we must have $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t_{2})\leq\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t_{1})$ for all $0\leq t_{1}<t_{2}<1$. Thus $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t)$ is non-increasing on $[0,1)$, and together with $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(1)=0$ this completes the proof. ∎
@@ Appendix B Numerical Solution to the Differential Equation in Theorem 1
We assume throughout that $\theta$ is left‑continuous, non‑increasing, and $\theta(s)\equiv 0$ for $s\in(\lambda_{2},1]$.
@@
Here, the last equality is obtained by exchanging the order of integration and using $\theta\equiv 0$ on $(\lambda_{2},1]$.
@@
where the left-hand side decreases from $\infty$ to 0 as $\theta_{m}$ increases from 0 to $\infty$. Thus, there exists a $\theta_{m}^{*}>0$ such that the equality holds. [^11] During experiments, we use binary search to locate $\theta_{m}^{*}$ with $100$ iterations so that the error can be ignored.
@@
By merging intervals where $\theta_{i}$ exceeds $1$, and by Lemma 26, we only need to check that Equation 27 holds for $\min\{\theta_{i},1\}$, where $i$ is the last index such that $\theta_{i}>1$. This follows immediately from the non-increasing property of the left-hand side. ∎
@@
If there exists an $\alpha$ such that the corresponding recursively constructed sequence $(\theta_{i}^{*})_{i=1}^{m}$ satisfies $\theta_{1}^{*}\geq 1$, then the step function $\theta^{*}$ with step values $(\min\{\theta_{i}^{*},1\})_{i=1}^{m}$ is a left-continuous non-increasing function satisfying Equation 23 and Equation 24.
@@
Intuitively, as $m\to\infty$, the step function approximation tightens. In experiments, we set $m=300$. For details, see the implementation of the backward recursion as well as the solved $\theta_{i}^{*}$’s in our code repository. [^12] https://github.com/billyldc/optimal-stopping-with-a-predicted-prior
Figure 3 illustrates the case where $\beta=0.01$ and $m=300$, which yields $\alpha\approx 0.6908$ with $\theta^{*}_{1}\approx 1.0165$. This implies that $\mathcal{A}(F,(\theta^{*})^{\nicefrac{{1}}{{n}}})$ with $\theta^{*}$ being the step function with step values $(\min\{\theta_{i}^{*},1\})_{i=1}^{300}$ is an $\alpha\approx 0.6908$-consistent and $\beta=0.01$-robust algorithm.
@@
Appendix C Missing Proofs for MaxProb Consistency
Lemma 30.
For any fixed $s\in[0,1)$, we have
$\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(s)~=~\frac{1}{1+\frac{\gamma}{(n-1)(1-s)}}+o\Big(\frac{1}{n}\Big)~,$
where $\gamma\approx 0.80435$ satisfies
$\sum_{k\geq 1}\frac{\gamma^{k}}{k!k}=1~.$
Proof.
Recall Equation 28:
$\int_{s}^{1}\frac{\big(\frac{1-t}{\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(s)}+t\big)^{n-1}-1}{1-t}\,\mathrm{d}t=1~.$ (28)
Fix any $s$, substitute $x_{n}=\frac{1}{\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(s)}-1$, the left-hand side becomes
$\int_{s}^{1}\frac{\left(1+x_{n}(1-t)\right)^{n-1}-1}{1-t}\,\mathrm{d}t~.$
Expanding the left-hand side with the binomial theorem, we have
$(n-1)(1-s)x_{n}+\binom{n-1}{2}\frac{(1-s)^{2}}{2}x_{n}^{2}+\cdots+\binom{n-1}{k}\frac{(1-s)^{k}}{k}x_{n}^{k}+\cdots+\frac{(1-s)^{n-1}}{n-1}x_{n}^{n-1}=1~.$
Intuitively, as $n\to\infty$, the coefficients $\frac{(n-2)(n-3)\cdots(n-k)}{(n-1)^{k-1}}\to 1$ for any fixed $k$. Thus, we expect that the value $c$ satisfies
$\sum_{k\geq 1}\frac{c^{k}}{k!k}=1~,$
which defines the constant $\gamma\approx 0.80435$ in the lemma.
Consider the difference:
It follows that the difference is bounded by
Finally, we conclude that $(n-1)(1-s)x_{n}\to\gamma$ as $n\to\infty$, which implies that $x_{n}=\frac{\gamma}{(n-1)(1-s)}+o(\tfrac{1}{n})$. This completes the proof by $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(s)=\frac{1}{1+x_{n}}=\frac{1}{1+\frac{\gamma}{(n-1)(1-s)}}+o(\tfrac{1}{n})$. ∎
Appendix D Missing Proofs for MaxProb Hardness
D.1 Minor-Oblivious Algorithms: Proof of Proposition 19
The key observation is that, when the current value is a record and equal across sequences, the conditional probability of stopping at time $t$ is invariant across all truncation levels $k\geq\ell$ for which the conditioning event has positive probability. This is formalized in the following lemma.
@@ D.1 Minor-Oblivious Algorithms: Proof of Proposition 19
That is, the probability of accepting a record $x_{t}=\ell$ equals the probability that $\mathcal{A}$ accepts at time $t$, conditioned on 1) $x_{t}=\ell$; 2) $x_{t}$ is the record; and 3) $\mathcal{A}$ rejects the first t1t-1 values. By Lemma 31, we may define this conditional probability with respect to prior $F_{K}$ without loss of generality.
@@
D.2 Relating Rejection and Acceptance Probabilities: Proof of Lemma 20